### v1

# Euclid Quick Data Release (Q1)

11institutetext: Dipartimento di Fisica e Scienze della Terra, Università degli Studi di Ferrara, Via Giuseppe Saragat 1, 44122 Ferrara, Italy 22institutetext: INAF-Osservatorio di Astrofisica e Scienza dello Spazio di Bologna, Via Piero Gobetti 93/3, 40129 Bologna, Italy 33institutetext: INAF-Osservatorio Astronomico di Capodimonte, Via Moiariello 16, 80131 Napoli, Italy 44institutetext: Dipartimento di Fisica ”Aldo Pontremoli”, Università degli Studi di Milano, Via Celoria 16, 20133 Milano, Italy 55institutetext: INFN-Sezione di Bologna, Viale Berti Pichat 6/2, 40127 Bologna, Italy 66institutetext: Instituto de Física de Cantabria, Edificio Juan Jordá, Avenida de los Castros, 39005 Santander, Spain 77institutetext: INAF-IASF Milano, Via Alfonso Corti 12, 20133 Milano, Italy 88institutetext: Universita di Salerno, Dipartimento di Fisica ”E.R. Caianiello”, Via Giovanni Paolo II 132, I-84084 Fisciano (SA), Italy 99institutetext: Institute of Cosmology and Gravitation, University of Portsmouth, Portsmouth PO1 3FX, UK 1010institutetext: INFN – Gruppo Collegato di Salerno - Sezione di Napoli, Dipartimento di Fisica ”E.R. Caianiello”, Universita di Salerno, via Giovanni Paolo II, 132 - I-84084 Fisciano (SA), Italy 1111institutetext: ESAC/ESA, Camino Bajo del Castillo, s/n., Urb. Villafranca del Castillo, 28692 Villanueva de la Cañada, Madrid, Spain 1212institutetext: INAF-Osservatorio Astronomico di Brera, Via Brera 28, 20122 Milano, Italy 1313institutetext: IFPU, Institute for Fundamental Physics of the Universe, via Beirut 2, 34151 Trieste, Italy 1414institutetext: INAF-Osservatorio Astronomico di Trieste, Via G. B. Tiepolo 11, 34143 Trieste, Italy 1515institutetext: INFN, Sezione di Trieste, Via Valerio 2, 34127 Trieste TS, Italy 1616institutetext: SISSA, International School for Advanced Studies, Via Bonomea 265, 34136 Trieste TS, Italy 1717institutetext: Dipartimento di Fisica e Astronomia, Università di Bologna, Via Gobetti 93/2, 40129 Bologna, Italy 1818institutetext: INAF-Osservatorio Astronomico di Padova, Via dell’Osservatorio 5, 35122 Padova, Italy 1919institutetext: Dipartimento di Fisica, Università di Genova, Via Dodecaneso 33, 16146, Genova, Italy 2020institutetext: INFN-Sezione di Genova, Via Dodecaneso 33, 16146, Genova, Italy 2121institutetext: Department of Physics ”E. Pancini”, University Federico II, Via Cinthia 6, 80126, Napoli, Italy 2222institutetext: Dipartimento di Fisica, Università degli Studi di Torino, Via P. Giuria 1, 10125 Torino, Italy 2323institutetext: INFN-Sezione di Torino, Via P. Giuria 1, 10125 Torino, Italy 2424institutetext: INAF-Osservatorio Astrofisico di Torino, Via Osservatorio 20, 10025 Pino Torinese (TO), Italy 2525institutetext: European Space Agency/ESTEC, Keplerlaan 1, 2201 AZ Noordwijk, The Netherlands 2626institutetext: Leiden Observatory, Leiden University, Einsteinweg 55, 2333 $C$ Leiden, The Netherlands 2727institutetext: Centro de Investigaciones Energéticas, Medioambientales y Tecnológicas (CIEMAT), Avenida Complutense 40, 28040 Madrid, Spain 2828institutetext: Port d’Informació Científica, Campus UAB, C. Albareda s/n, 08193 Bellaterra (Barcelona), Spain 2929institutetext: INAF-Osservatorio Astronomico di Roma, Via Frascati 33, 00078 Monteporzio Catone, Italy 3030institutetext: INFN section of Naples, Via Cinthia 6, 80126, Napoli, Italy 3131institutetext: Dipartimento di Fisica e Astronomia ”Augusto Righi” - Alma Mater Studiorum Università di Bologna, Viale Berti Pichat 6/2, 40127 Bologna, Italy 3232institutetext: Instituto de Astrofísica de Canarias, Vía Láctea, 38205 La Laguna, Tenerife, Spain 3333institutetext: Institute for Astronomy, University of Edinburgh, Royal Observatory, Blackford Hill, Edinburgh EH9 3HJ, UK 3434institutetext: European Space Agency/ESRIN, Largo Galileo Galilei 1, 00044 Frascati, Roma, Italy 3535institutetext: Université Claude Bernard Lyon 1, CNRS/IN2P3, IP2I Lyon, UMR 5822, Villeurbanne, F-69100, France 3636institutetext: Aix-Marseille Université, CNRS, CNES, LAM, Marseille, France 3737institutetext: Institut de Ciències del Cosmos (ICCUB), Universitat de Barcelona (IEEC-UB), Martí i Franquès 1, 08028 Barcelona, Spain 3838institutetext: Institució Catalana de Recerca i Estudis Avançats (ICREA), Passeig de Lluís Companys 23, 08010 Barcelona, Spain 3939institutetext: Institut de Ciencies de l’Espai (IEEC-CSIC), Campus UAB, Carrer de Can Magrans, s/n Cerdanyola del Vallés, 08193 Barcelona, Spain 4040institutetext: UCB Lyon 1, CNRS/IN2P3, IUF, IP2I Lyon, 4 rue Enrico Fermi, 69622 Villeurbanne, France 4141institutetext: Mullard Space Science Laboratory, University College London, Holmbury St Mary, Dorking, Surrey RH5 6NT, UK 4242institutetext: Departamento de Física, Faculdade de Ciências, Universidade de Lisboa, Edifício C8, Campo Grande, PT1749-016 Lisboa, Portugal 4343institutetext: Instituto de Astrofísica e Ciências do Espaço, Faculdade de Ciências, Universidade de Lisboa, Campo Grande, 1749-016 Lisboa, Portugal 4444institutetext: Department of Astronomy, University of Geneva, ch. d’Ecogia 16, 1290 Versoix, Switzerland 4545institutetext: Université Paris-Saclay, CNRS, Institut d’astrophysique spatiale, 91405, Orsay, France 4646institutetext: INFN-Padova, Via Marzolo 8, 35131 Padova, Italy 4747institutetext: Aix-Marseille Université, CNRS/IN2P3, CPPM, Marseille, France 4848institutetext: Max Planck Institute for Extraterrestrial Physics, Giessenbachstr. 1, 85748 Garching, Germany 4949institutetext: Universitäts-Sternwarte München, Fakultät für Physik, Ludwig-Maximilians-Universität München, Scheinerstrasse 1, 81679 München, Germany 5050institutetext: INAF-Istituto di Astrofisica e Planetologia Spaziali, via del Fosso del Cavaliere, 100, 00100 Roma, Italy 5151institutetext: Space Science Data Center, Italian Space Agency, via del Politecnico snc, 00133 Roma, Italy 5252institutetext: INFN-Bologna, Via Irnerio 46, 40126 Bologna, Italy 5353institutetext: University Observatory, LMU Faculty of Physics, Scheinerstrasse 1, 81679 Munich, Germany 5454institutetext: INFN-Sezione di Milano, Via Celoria 16, 20133 Milano, Italy 5555institutetext: Institute of Theoretical Astrophysics, University of Oslo, P.O. Box 1029 Blindern, 0315 Oslo, Norway 5656institutetext: Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Drive, Pasadena, CA, 91109, USA 5757institutetext: Department of Physics, Lancaster University, Lancaster, LA1 4YB, UK 5858institutetext: Felix Hormuth Engineering, Goethestr. 17, 69181 Leimen, Germany 5959institutetext: Technical University of Denmark, Elektrovej 327, 2800 Kgs. Lyngby, Denmark 6060institutetext: Cosmic Dawn Center (DAWN), Denmark 6161institutetext: Max-Planck-Institut für Astronomie, Königstuhl 17, 69117 Heidelberg, Germany 6262institutetext: NASA Goddard Space Flight Center, Greenbelt, MD 20771, USA 6363institutetext: Department of Physics and Astronomy, University College London, Gower Street, London WC1E 6BT, UK 6464institutetext: Department of Physics and Helsinki Institute of Physics, Gustaf Hällströmin katu 2, University of Helsinki, 00014 Helsinki, Finland 6565institutetext: Université Paris-Saclay, Université Paris Cité, CEA, CNRS, AIM, 91191, Gif-sur-Yvette, France 6666institutetext: Université de Genève, Département de Physique Théorique and Centre for Astroparticle Physics, 24 quai Ernest-Ansermet, CH-1211 Genève 4, Switzerland 6767institutetext: Department of Physics, P.O. Box 64, University of Helsinki, 00014 Helsinki, Finland 6868institutetext: Helsinki Institute of Physics, Gustaf Hällströmin katu 2, University of Helsinki, 00014 Helsinki, Finland 6969institutetext: Kapteyn Astronomical Institute, University of Groningen, PO Box 800, 9700 AV Groningen, The Netherlands 7070institutetext: Laboratoire d’etude de l’Univers et des phenomenes eXtremes, Observatoire de Paris, Université PSL, Sorbonne Université, CNRS, 92190 Meudon, France 7171institutetext: SKAO, Jodrell Bank, Lower Withington, Macclesfield SK11 9FT, United Kingdom 7272institutetext: Centre de Calcul de l’IN2P3/CNRS, 21 avenue Pierre de Coubertin 69627 Villeurbanne Cedex, France 7373institutetext: Universität Bonn, Argelander-Institut für Astronomie, Auf dem Hügel 71, 53121 Bonn, Germany 7474institutetext: INFN-Sezione di Roma, Piazzale Aldo Moro, 2 - c/o Dipartimento di Fisica, Edificio G. Marconi, 00185 Roma, Italy 7575institutetext: Dipartimento di Fisica e Astronomia ”Augusto Righi” - Alma Mater Studiorum Università di Bologna, via Piero Gobetti 93/2, 40129 Bologna, Italy 7676institutetext: Department of Physics, Institute for Computational Cosmology, Durham University, South Road, Durham, DH1 3LE, UK 7777institutetext: Université Paris Cité, CNRS, Astroparticule et Cosmologie, 75013 Paris, France 7878institutetext: CNRS-UCB International Research Laboratory, Centre Pierre Binétruy, IRL2007, CPB-IN2P3, Berkeley, USA 7979institutetext: University of Applied Sciences and Arts of Northwestern Switzerland, School of Engineering, 5210 Windisch, Switzerland 8080institutetext: Institut d’Astrophysique de Paris, 98bis Boulevard Arago, 75014, Paris, France 8181institutetext: Institut d’Astrophysique de Paris, UMR 7095, CNRS, and Sorbonne Université, 98 bis boulevard Arago, 75014 Paris, France 8282institutetext: Institute of Physics, Laboratory of Astrophysics, Ecole Polytechnique Fédérale de Lausanne (EPFL), Observatoire de Sauverny, 1290 Versoix, Switzerland 8383institutetext: Telespazio UK S.L. for European Space Agency (ESA), Camino bajo del Castillo, s/n, Urbanizacion Villafranca del Castillo, Villanueva de la Cañada, 28692 Madrid, Spain 8484institutetext: Institut de Física d’Altes Energies (IFAE), The Barcelona Institute of Science and Technology, Campus UAB, 08193 Bellaterra (Barcelona), Spain 8585institutetext: DARK, Niels Bohr Institute, University of Copenhagen, Jagtvej 155, 2200 Copenhagen, Denmark 8686institutetext: Waterloo Centre for Astrophysics, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada 8787institutetext: Department of Physics and Astronomy, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada 8888institutetext: Perimeter Institute for Theoretical Physics, Waterloo, Ontario N2L 2Y5, Canada 8989institutetext: Centre National d’Etudes Spatiales – Centre spatial de Toulouse, 18 avenue Edouard Belin, 31401 Toulouse Cedex 9, France 9090institutetext: Institute of Space Science, Str. Atomistilor, nr. 409 Măgurele, Ilfov, 077125, Romania 9191institutetext: Dipartimento di Fisica e Astronomia ”G. Galilei”, Università di Padova, Via Marzolo 8, 35131 Padova, Italy 9292institutetext: Institut für Theoretische Physik, University of Heidelberg, Philosophenweg 16, 69120 Heidelberg, Germany 9393institutetext: Institut de Recherche en Astrophysique et Planétologie (IRAP), Université de Toulouse, CNRS, UPS, CNES, 14 Av. Edouard Belin, 31400 Toulouse, France 9494institutetext: Université St Joseph; Faculty of Sciences, Beirut, Lebanon 9595institutetext: Departamento de Física, FCFM, Universidad de Chile, Blanco Encalada 2008, Santiago, Chile 9696institutetext: Universität Innsbruck, Institut für Astro- und Teilchenphysik, Technikerstr. 25/8, 6020 Innsbruck, Austria 9797institutetext: Institut d’Estudis Espacials de Catalunya (IEEC), Edifici RDIT, Campus UPC, 08860 Castelldefels, Barcelona, Spain 9898institutetext: Satlantis, University Science Park, Sede Bld 48940, Leioa-Bilbao, Spain 9999institutetext: Institute of Space Sciences (ICE, CSIC), Campus UAB, Carrer de Can Magrans, s/n, 08193 Barcelona, Spain 100100institutetext: Instituto de Astrofísica e Ciências do Espaço, Faculdade de Ciências, Universidade de Lisboa, Tapada da Ajuda, 1349-018 Lisboa, Portugal 101101institutetext: Cosmic Dawn Center (DAWN) 102102institutetext: Niels Bohr Institute, University of Copenhagen, Jagtvej 128, 2200 Copenhagen, Denmark 103103institutetext: Universidad Politécnica de Cartagena, Departamento de Electrónica y Tecnología de Computadoras, Plaza del Hospital 1, 30202 Cartagena, Spain 104104institutetext: Infrared Processing and Analysis Center, California Institute of Technology, Pasadena, CA 91125, USA 105105institutetext: Istituto Nazionale di Fisica Nucleare, Sezione di Ferrara, Via Giuseppe Saragat 1, 44122 Ferrara, Italy 106106institutetext: INAF, Istituto di Radioastronomia, Via Piero Gobetti 101, 40129 Bologna, Italy 107107institutetext: Astronomical Observatory of the Autonomous Region of the Aosta Valley (OAVdA), Loc. Lignan 39, I-11020, Nus (Aosta Valley), Italy 108108institutetext: Université Côte d’Azur, Observatoire de la Côte d’Azur, CNRS, Laboratoire Lagrange, Bd de l’Observatoire, CS 34229, 06304 Nice cedex 4, France 109109institutetext: Aurora Technology for European Space Agency (ESA), Camino bajo del Castillo, s/n, Urbanizacion Villafranca del Castillo, Villanueva de la Cañada, 28692 Madrid, Spain 110110institutetext: ICL, Junia, Université Catholique de Lille, LITL, 59000 Lille, France 111111institutetext: ICSC - Centro Nazionale di Ricerca in High Performance Computing, Big Data e Quantum Computing, Via Magnanelli 2, Bologna, Italy 112112institutetext: Instituto de Física Teórica UAM-CSIC, Campus de Cantoblanco, 28049 Madrid, Spain 113113institutetext: CERCA/ISO, Department of Physics, Case Western Reserve University, 10900 Euclid Avenue, Cleveland, OH 44106, USA 114114institutetext: Technical University of Munich, TUM School of Natural Sciences, Physics Department, James-Franck-Str. 1, 85748 Garching, Germany 115115institutetext: Max-Planck-Institut für Astrophysik, Karl-Schwarzschild-Str. 1, 85748 Garching, Germany 116116institutetext: Laboratoire Univers et Théorie, Observatoire de Paris, Université PSL, Université Paris Cité, CNRS, 92190 Meudon, France 117117institutetext: Departamento de Física Fundamental. Universidad de Salamanca. Plaza de la Merced s/n. 37008 Salamanca, Spain 118118institutetext: IRFU, CEA, Université Paris-Saclay 91191 Gif-sur-Yvette Cedex, France 119119institutetext: Université de Strasbourg, CNRS, Observatoire astronomique de Strasbourg, UMR 7550, 67000 Strasbourg, France 120120institutetext: Center for Data-Driven Discovery, Kavli IPMU (WPI), UTIAS, The University of Tokyo, Kashiwa, Chiba 277-8583, Japan 121121institutetext: Ludwig-Maximilians-University, Schellingstrasse 4, 80799 Munich, Germany 122122institutetext: Max-Planck-Institut für Physik, Boltzmannstr. 8, 85748 Garching, Germany 123123institutetext: Dipartimento di Fisica - Sezione di Astronomia, Università di Trieste, Via Tiepolo 11, 34131 Trieste, Italy 124124institutetext: Jodrell Bank Centre for Astrophysics, Department of Physics and Astronomy, University of Manchester, Oxford Road, Manchester M13 9PL, UK 125125institutetext: California Institute of Technology, 1200 E California Blvd, Pasadena, CA 91125, USA 126126institutetext: SCITAS, Ecole Polytechnique Fédérale de Lausanne (EPFL), 1015 Lausanne, Switzerland 127127institutetext: Department of Physics & Astronomy, University of California Irvine, Irvine CA 92697, USA 128128institutetext: Department of Mathematics and Physics E. De Giorgi, University of Salento, Via per Arnesano, CP-I93, 73100, Lecce, Italy 129129institutetext: INFN, Sezione di Lecce, Via per Arnesano, CP-193, 73100, Lecce, Italy 130130institutetext: INAF-Sezione di Lecce, c/o Dipartimento Matematica e Fisica, Via per Arnesano, 73100, Lecce, Italy 131131institutetext: Departamento Física Aplicada, Universidad Politécnica de Cartagena, Campus Muralla del Mar, 30202 Cartagena, Murcia, Spain 132132institutetext: Observatorio Nacional, Rua General Jose Cristino, 77-Bairro Imperial de Sao Cristovao, Rio de Janeiro, 20921-400, Brazil 133133institutetext: CEA Saclay, DFR/IRFU, Service d’Astrophysique, Bat. 709, 91191 Gif-sur-Yvette, France 134134institutetext: Department of Astronomy, University of Florida, Bryant Space Science Center, Gainesville, FL 32611, USA 135135institutetext: Department of Computer Science, Aalto University, PO Box 15400, Espoo, FI-$0$ 076, Finland 136136institutetext: Instituto de Astrofísica de Canarias, c/ Via Lactea s/n, La Laguna 38200, Spain. Departamento de Astrofísica de la Universidad de La Laguna, Avda. Francisco Sanchez, La Laguna, 38200, Spain 137137institutetext: Caltech/IPAC, 1200 E. California Blvd., Pasadena, CA 91125, USA 138138institutetext: Ruhr University Bochum, Faculty of Physics and Astronomy, Astronomical Institute (AIRUB), German Centre for Cosmological Lensing (GCCL), 44780 Bochum, Germany 139139institutetext: Department of Physics and Astronomy, Vesilinnantie 5, University of Turku, 20014 Turku, Finland 140140institutetext: Serco for European Space Agency (ESA), Camino bajo del Castillo, s/n, Urbanizacion Villafranca del Castillo, Villanueva de la Cañada, 28692 Madrid, Spain 141141institutetext: ARC Centre of Excellence for Dark Matter Particle Physics, Melbourne, Australia 142142institutetext: Centre for Astrophysics & Supercomputing, Swinburne University of Technology, Hawthorn, Victoria 3122, Australia 143143institutetext: Department of Physics and Astronomy, University of the Western Cape, Bellville, Cape Town, 7535, South Africa 144144institutetext: Department of Physics, Oxford University, Keble Road, Oxford OX1 3RH, UK 145145institutetext: DAMTP, Centre for Mathematical Sciences, Wilberforce Road, Cambridge CB3 0WA, UK 146146institutetext: Kavli Institute for Cosmology Cambridge, Madingley Road, Cambridge, CB3 0HA, UK 147147institutetext: Department of Astrophysics, University of Zurich, Winterthurerstrasse 190, 8057 Zurich, Switzerland 148148institutetext: Department of Physics, Centre for Extragalactic Astronomy, Durham University, South Road, Durham, DH1 3LE, UK 149149institutetext: Institute for Theoretical Particle Physics and Cosmology (TTK), RWTH Aachen University, 52056 Aachen, Germany 150150institutetext: Oskar Klein Centre for Cosmoparticle Physics, Department of Physics, Stockholm University, Stockholm, SE-106 91, Sweden 151151institutetext: Astrophysics Group, Blackett Laboratory, Imperial College London, London SW7 2AZ, UK 152152institutetext: Univ. Grenoble Alpes, CNRS, Grenoble INP, LPSC-IN2P3, 53, Avenue des Martyrs, 38000, Grenoble, France 153153institutetext: INAF-Osservatorio Astrofisico di Arcetri, Largo E. Fermi 5, 50125, Firenze, Italy 154154institutetext: Dipartimento di Fisica, Sapienza Università di Roma, Piazzale Aldo Moro 2, 00185 Roma, Italy 155155institutetext: Centro de Astrofísica da Universidade do Porto, Rua das Estrelas, 4150-762 Porto, Portugal 156156institutetext: Instituto de Astrofísica e Ciências do Espaço, Universidade do Porto, CAUP, Rua das Estrelas, PT4150-762 Porto, Portugal 157157institutetext: HE Space for European Space Agency (ESA), Camino bajo del Castillo, s/n, Urbanizacion Villafranca del Castillo, Villanueva de la Cañada, 28692 Madrid, Spain 158158institutetext: INAF - Osservatorio Astronomico d’Abruzzo, Via Maggini, 64100, Teramo, Italy 159159institutetext: Theoretical astrophysics, Department of Physics and Astronomy, Uppsala University, Box 516, 751 37 Uppsala, Sweden 160160institutetext: Institute for Astronomy, University of Hawaii, 2680 Woodlawn Drive, Honolulu, HI 96822, USA 161161institutetext: Mathematical Institute, University of Leiden, Einsteinweg 55, 2333 CA Leiden, The Netherlands 162162institutetext: Institute of Astronomy, University of Cambridge, Madingley Road, Cambridge CB3 0HA, UK 163163institutetext: Univ. Lille, CNRS, Centrale Lille, UMR 9189 CRIStAL, 59000 Lille, France 164164institutetext: Department of Astrophysical Sciences, Peyton Hall, Princeton University, Princeton, NJ 08544, USA 165165institutetext: Space physics and astronomy research unit, University of Oulu, Pentti Kaiteran katu 1, FI-90014 Oulu, Finland 166166institutetext: Center for Computational Astrophysics, Flatiron Institute, 162 5th Avenue, 10010, New York, NY, USA
Euclid Quick Data Release (Q1)
Searching for giant gravitational arcs in galaxy clusters with mask region-based convolutional neural networks
Euclid Collaboration: L. Bazzanini    G. Angora    P. Bergamini    M. Meneghetti    P. Rosati    A. Acebron    C. Grillo    M. Lombardi    R. Ratta    M. Fogliardi    G. Di Rosa    D. Abriola    M. D’Addona    G. Granata    L. Leuzzi    A. Mercurio    S. Schuldt    E. Vanzella    C. Tortora    B. Altieri    S. Andreon    N. Auricchio    C. Baccigalupi    M. Baldi    A. Balestra    S. Bardelli    P. Battaglia    A. Biviano    E. Branchini    M. Brescia    S. Camera    G. Cañas-Herrera    V. Capobianco    C. Carbone    J. Carretero    M. Castellano    G. Castignani    S. Cavuoti    A. Cimatti    C. Colodro-Conde    G. Congedo    L. Conversi    Y. Copin    A. Costille    F. Courbin    H. M. Courtois    M. Cropper    A. Da Silva    H. Degaudenzi    G. De Lucia    H. Dole    F. Dubath    C. A. J. Duncan    X. Dupac    S. Dusini    S. Escoffier    M. Fabricius    M. Farina    R. Farinelli    F. Faustini    S. Ferriol    F. Finelli    M. Frailis    E. Franceschi    M. Fumana    S. Galeotta    K. George    W. Gillard    B. Gillis    C. Giocoli    J. Gracia-Carpio    A. Grazian    F. Grupp    L. Guzzo    S. V. H. Haugan    J. Hoar    W. Holmes    I. M. Hook    F. Hormuth    A. Hornstrup    K. Jahnke    M. Jhabvala    B. Joachimi    E. Keihänen    S. Kermiche    A. Kiessling    M. Kilbinger    B. Kubik    M. Kunz    H. Kurki-Suonio    R. Laureijs    A. M. C. Le Brun    D. Le Mignant    S. Ligori    P. B. Lilje    V. Lindholm    I. Lloro    G. Mainetti    D. Maino    E. Maiorano    O. Mansutti    O. Marggraf    M. Martinelli    N. Martinet    F. Marulli    R. J. Massey    E. Medinaceli    S. Mei    M. Melchior    Y. Mellier    E. Merlin    G. Meylan    A. Mora    M. Moresco    L. Moscardini    C. Neissner    S.-M. Niemi    C. Padilla    S. Paltani    F. Pasian    K. Pedersen    W. J. Percival    V. Pettorino    S. Pires    G. Polenta    M. Poncet    L. A. Popa    L. Pozzetti    F. Raison    A. Renzi    J. Rhodes    G. Riccio    E. Romelli    M. Roncarelli    R. Saglia    Z. Sakr    A. G. Sánchez    D. Sapone    B. Sartoris    P. Schneider    T. Schrabback    A. Secroun    G. Seidel    S. Serrano    P. Simon    C. Sirignano    G. Sirri    L. Stanco    J. Steinwagner    P. Tallada-Crespí    A. N. Taylor    I. Tereno    N. Tessore    S. Toft    R. Toledo-Moreo    F. Torradeflot    I. Tutusaus    E. A. Valentijn    L. Valenziano    J. Valiviita    T. Vassallo    G. Verdoes Kleijn    A. Veropalumbo    Y. Wang    J. Weller    A. Zacchei    G. Zamorani    E. Zucca    M. Ballardini    M. Bolzonella    E. Bozzo    C. Burigana    R. Cabanac    M. Calabrese    A. Cappi    D. Di Ferdinando    J. A. Escartin Vigo    W. G. Hartley    J. Martín-Fleitas    S. Matthew    N. Mauri    R. B. Metcalf    A. Pezzotta    M. Pöntinen    I. Risso    V. Scottez    M. Sereno    M. Tenti    M. Viel    M. Wiesmann    Y. Akrami    I. T. Andika    S. Anselmi    M. Archidiacono    F. Atrio-Barandela    E. Aubourg    D. Bertacca    M. Bethermin    A. Blanchard    L. Blot    H. Böhringer    M. Bonici    S. Borgani    M. L. Brown    S. Bruton    A. Calabro    B. Camacho Quevedo    F. Caro    C. S. Carvalho    T. Castro    B. Clément    F. Cogato    S. Conseil    A. R. Cooray    O. Cucciati    S. Davini    F. De Paolis    G. Desprez    A. Díaz-Sánchez    J. J. Diaz    S. Di Domizio    J. M. Diego    P. Dimauro    P.-A. Duc    M. Y. Elkhashab    A. Enia    Y. Fang    A. Finoguenov    A. Fontana    A. Franco    K. Ganga    J. García-Bellido    T. Gasparetto    V. Gautard    R. Gavazzi    E. Gaztanaga    F. Giacomini    F. Gianotti    A. H. Gonzalez    G. Gozaliasl    M. Guidi    C. M. Gutierrez    S. Hemmati    H. Hildebrandt    J. Hjorth    J. J. E. Kajava    Y. Kang    V. Kansal    D. Karagiannis    K. Kiiveri    J. Kim    C. C. Kirkpatrick    S. Kruk    J. Le Graet    L. Legrand    M. Lembo    F. Lepori    G. Leroy    G. F. Lesci    J. Lesgourgues    T. I. Liaudat    S. J. Liu    A. Loureiro    J. Macias-Perez    M. Magliocchetti    F. Mannucci    R. Maoli    C. J. A. P. Martins    L. Maurin    C. J. R. McPartland    M. Miluzio    P. Monaco    C. Moretti    G. Morgante    C. Murray    K. Naidoo    A. Navarro-Alsina    S. Nesseris    D. Paoletti    F. Passalacqua    K. Paterson    A. Pisani    D. Potter    S. Quai    M. Radovich    P.-F. Rocci    S. Sacquegna    M. Sahlén    D. B. Sanders    E. Sarpa    A. Schneider    D. Sciotti    E. Sellentin    L. C. Smith    J. G. Sorce    K. Tanidis    C. Tao    G. Testera    R. Teyssier    S. Tosi    A. Troja    M. Tucci    C. Valieri    A. Venhola    D. Vergani    G. Verza    P. Vielzeuf    N. A. Walton
(November 4, 2025)
Strong gravitational lensing (SL) by galaxy clusters is a powerful probe of their inner mass distribution and a key test bed for cosmological models. However, the detection of SL events in wide-field surveys such as Euclid requires robust, automated methods capable of handling the immense data volume generated. In this work, we present an advanced deep learning (DL) framework based on mask region-based convolutional neural networks (Mask R-CNNs), designed to autonomously detect and segment bright, strongly-lensed arcs in Euclid’s multi-band imaging of galaxy clusters. The model is trained on a realistic simulated data set of cluster-scale SL events, constructed by injecting mock background sources into Euclidised Hubble Space Telescope images of $10$ massive lensing clusters, exploiting their high-precision mass models constructed with extensive spectroscopic data. The network is trained and validated on over $4500$ simulated images, and tested on an independent set of $500$ simulations, as well as real Euclid Quick Data Release (Q1) observations. The trained network achieves high performance in identifying gravitational arcs in the test set, with a precision and recall of $76\%$ and $58\%$, respectively, processing $2^{\prime}\times 2^{\prime}$ images in a fraction of a second. When applied to a sample of visually confirmed Euclid Q1 cluster-scale lenses, our model recovers $\approx 66\%$ of gravitational arcs above the area threshold used during training. While the model shows promising results, limitations include the production of some false positives and challenges in detecting smaller, fainter arcs. Our results demonstrate the potential of advanced DL computer vision techniques for efficient and scalable arc detection, enabling the automated analysis of SL systems in current and future wide-field surveys. The code, ARTEMIDE, is open source and will be available at github.com/LBasz/ARTEMIDE.
Key Words.:
Gravitational lensing: strong – Galaxies: clusters: general – Techniques: image processing
1 Introduction
Strong gravitational lensing (SL) serves as a valuable tool for probing the total mass distribution in galaxies and galaxy clusters, as well as testing cosmological total models (Meneghetti 2021). Over the past decades, SL has been utilised in various contexts: to study galaxy structures and their evolution (Treu & Koopmans 2002; Auger et al. 2010; Sonnenfeld et al. 2013); to estimate the Hubble constant ($H_{0}$) through time-delay measurements (Suyu et al. 2017; Grillo et al. 2018; Millon et al. 2020; Moresco et al. 2022; Grillo et al. 2024; TDCOSMO Collaboration et al. 2025); to constrain the dark energy (DE) equation of state (Jullo et al. 2010; Cao et al. 2012; Collett & Auger 2014; Caminha et al. 2022); and to estimate the dark matter (DM) fraction in massive early-type galaxies (Auger et al. 2010; Tortora et al. 2010; Sonnenfeld et al. 2015). On galaxy cluster scales, SL is instrumental in modelling their inner total mass distribution, leveraging multiple images of lensed background sources (Caminha et al. 2017, 2019; Acebron et al. 2018; Lagattuta et al. 2019; Bergamini et al. 2019, 2021; Lagattuta et al. 2022; Bergamini et al. 2023a, b). Furthermore, the magnification effect of SL allows galaxy clusters to act as cosmic telescopes, enabling the study of faint, high-redshift sources that would otherwise be unobservable (Swinbank et al. 2009; Richard et al. 2011; Vanzella et al. 2020, 2021).
Traditionally, visual inspection has been the primary method for confirming lens candidates, typically following an initial pre-selection based on spectroscopic, photometric, or morphological criteria (e.g. Le Fèvre & Hammer 1988; Jackson 2008; Sygnet et al. 2010; Pawase et al. 2014; Euclid Collaboration: Bergamini et al. 2025a). However, current and next-generation astronomical surveys, such as Euclid (Euclid Collaboration: Mellier et al. 2025), the Nancy Grace Roman Space Telescope (Green et al. 2012), and the Vera Rubin Observatory Legacy Survey of Space and Time (Ivezić et al. 2019), are expected to produce unprecedented volumes of imaging data over the coming decade. For example, Euclid alone is predicted to yield approximately $10^{5}$ galaxy-galaxy SL events (Collett 2015; Acevedo Barroso et al. 2025; Euclid Collaboration: Walmsley et al. 2025), as well as around $5000$ strongly lensing clusters (Euclid Collaboration: Bergamini et al. 2025a) out of the $10^{5}$ galaxy clusters up to $z=2$ it is expected to detect (Euclid Collaboration: Adam et al. 2019).
The visual inspection of Euclid Q1 (Euclid Quick Release Q1 2025) cluster lens candidates carried out in Euclid Collaboration: Bergamini et al. (2025a) involved the coordinated effort of around $40$ expert astronomers to assess approximately $1300$ candidate clusters within large $4^{\prime}\times 4^{\prime}$ cutouts. This task required several weeks of distributed work, underscoring the resource-intensive nature of such visual inspection. While effective, this approach is not sustainable for future data releases, which will contain orders of magnitude more data; our goal with this work is precisely to address this challenge. This underscores the need for more efficient, scalable, and automated approaches to analyse the vast amounts of data produced by these surveys. In the last decade, several alternative methods have been developed for detecting SL events, particularly on galaxy-scale. These methods range from semi-automated algorithms designed to detect arc and ring-like structures (e.g. More et al. 2012; Gavazzi et al. 2014; Sonnenfeld et al. 2018) to crowd-sourced science initiatives (e.g. Marshall et al. 2016; Sonnenfeld et al. 2020). Within this framework, machine learning (ML) and deep learning (DL) techniques have emerged as the most promising tools for identifying SL events (e.g. Metcalf et al. 2019).
Astronomy has benefited from applying ML and DL to both simulated and real survey data. These applications include, for instance, star detection (He et al. 2023), quasar and galaxy classification (Bailer-Jones et al. 2019; Clarke et al. 2020), photometric redshift estimation (D’Isanto & Polsterer 2018; Schuldt et al. 2021), cluster member classification (Angora et al. 2020), optimisation of stochastic models for the generation of gamma-ray burst light curves (Bazzanini et al. 2024); for an in-depth review we refer to Huertas-Company & Lanusse (2023). Due to their efficiency and high performance in image pattern recognition, supervised DL techniques, particularly convolutional neural networks (CNNs; LeCun et al. 1998, 2015), are becoming increasingly crucial for analysing astrophysical data sets. CNNs indeed have emerged as a powerful tool for identifying SL systems in imaging surveys (e.g. Petrillo et al. 2017; Cañameras et al. 2020; Huang et al. 2020; Jia et al. 2023; Angora et al. 2023; Euclid Collaboration: Walmsley et al. 2025). Once trained, these networks can process single images in a fraction of a second.
The field of computer vision has recently greatly advanced in areas such as object detection, classification, and semantic segmentation. The mask region-based convolutional neural network framework (Mask R-CNN; He et al. 2017) has set the benchmark for DL models in instance segmentation. Mask R-CNNs have been applied in astrophysics for detection, deblending, and classification of astronomical sources (Burke et al. 2019; Merz et al. 2024), for detection and morphological classification of galaxies (Farias et al. 2020), for detecting ‘ghosting artefacts’ (Tanoglidis et al. 2022), for detecting magnetic bright points in the solar photosphere (Yang et al. 2019, 2023), and for the detection and morphological classification of extended radio sources (Lao et al. 2023).
Deep learning methods, however, require training on suitable simulated data sets due to the limited number of confirmed SL systems, especially in galaxy clusters. To this end, considerable efforts have been made in recent years to simulate SL events akin to those observed by both ongoing and upcoming surveys. These simulations typically consist of generating mock images of SL events by superimposing simulated lensed sources onto foreground galaxies using various techniques (Meneghetti et al. 2008, 2010; Metcalf et al. 2019; Euclid Collaboration: Leuzzi et al. 2024). These sources are then combined with real or synthetic images using ray-tracing techniques, e.g. GLAMER (Metcalf & Petkova 2014; Petkova et al. 2014) and GRAVLENS (Keeton 2001).
In this work, we focus on detecting strongly lensed bright arcs on galaxy cluster scales, rather than galaxy-scale lensing events. Compared to galaxy-scale SL, typically searched for in small cutouts (e.g. $10^{\prime\prime}\times 10^{\prime\prime}$, Euclid Collaboration: Walmsley et al. 2025), cluster-scale arc detection presents additional challenges. The larger field of view used in our case ($2^{\prime}\times 2^{\prime}$) contains hundreds of sources per image, increasing the likelihood of source confusion and false positives. Furthermore, the cluster environment is more complex, featuring numerous bright and extended foreground galaxies and diverse arc morphologies caused by the complex multi-component mass distribution. These factors make automated detection particularly demanding.
The challenges of gravitational arc detection in galaxy clusters are well-suited for the Mask R-CNN framework; this advanced neural network (NN) indeed offers key advantages over standard CNNs. In particular, it can process full-size images of galaxy clusters thanks to its ability to handle inputs of variable size, without requiring resizing to a fixed input dimension, thereby preserving morphological features and spatial information. Moreover, it enables simultaneous detection, classification, and segmentation of multiple objects in a single image at different scales, allowing it to distinguish individual arc instances within crowded fields. These capabilities make Mask R-CNNs an effective choice for addressing the complexity of the cluster-scale lensing regime, positioning this work as an innovative application of advanced computer vision techniques to SL in wide-field surveys.
We train, validate, and test the performance of a Mask R-CNN using simulated cluster-scale SL events in Euclid simulated images (Euclid Collaboration: Bergamini et al. 2025b). The mock Euclid imaging data is generated from Hubble Space Telescope (HST) observations. To build the training set, we simulate thousands of galaxy cluster strong lensing (GCSL) arcs by taking advantage of high-precision cluster lens models constructed by Caminha et al. (2017, 2019); Bergamini et al. (2019, 2021, 2023a, 2023b) using the public software lenstool (Kneib et al. 1996; Jullo et al. 2007; Jullo & Kneib 2009). We then further test the NN by applying it to real Euclid images of lensing clusters found in Euclid Collaboration: Bergamini et al. (2025a), comparing the results of the NN inference with the lensing events found with a visual inspection. After training, this method is extremely efficient, detecting gravitational arcs in Euclid in a $3\times 1199\times 1199\,\mathrm{pixels}$ image in a fraction of a second using a single NVIDIA Quadro RTX 6000 GPU. Our code, ARTEMIDE (ARcs in clusTErs using Mask r-cnn IDEntifier), is free and open source, and available in our GitHub repository [^1] https://github.com/LBasz/ARTEMIDE.
This paper is organised as follows. In Sect. 2 we describe how the training set is produced via SL simulations. In Sect. 3, we introduce the Mask R-CNN framework, describing in detail the architecture of our implementation, and the training setup. In Sect. 4, we present the results of the trained network, we evaluate its performance on the test set, and we further test it on real Euclid Q1 images. Finally, in Sect. 5 we summarise, discuss the results, and conclude.
Throughout this paper, we assume a flat $\Lambda\mathrm{CDM}$ cosmology with $\Omega_{\Lambda}=0.7$, $\Omega_{\mathrm{m}}=0.3$, and $H_{0}=70\,\mathrm{km}\,\mathrm{s}^{-1}\,\mathrm{Mpc}^{-1}$. Magnitudes are reported in the AB system (Oke & Gunn 1983), unless otherwise stated.
2 Data set
In this section, we outline the procedure for generating the training set. We support multi-band Flexible Image Transport System files (FITS; Pence et al. 2010) as image input for the Mask R-CNN. In order to train a NN with a supervised learning approach, we typically need a large training set of lensing and non-lensing examples. However, the number of real cluster-scale gravitational arcs is limited to a small sample; therefore, to produce the ‘positive’ class examples we have to resort to SL simulations.
Since the most accurate lens models of galaxy clusters are available for objects observed by the HST, to generate the training set we had to degrade the HST images, in order to reproduce the observational parameters of the Euclid survey. We perform this task starting from the Python code HST2EUCLID (Euclid Collaboration: Bergamini et al. 2025b), a tool for converting HST observations into Euclid-like imaging data. With this software, we created the simulated data set of mock Euclid images, starting from $10$ HST observations of galaxy clusters (see Table 1) observed in the Cluster Lensing and Supernova Survey with Hubble (CLASH [^2] https://archive.stsci.edu/prepds/clash/; Postman et al. 2012) and the Hubble Frontier Fields (HFF [^3] https://archive.stsci.edu/prepds/frontier/; Lotz et al. 2014, 2017) programmes. Since the mock images are based on real observations, they inherently capture all the complexities present in the observed galaxy clusters.
Table 1: Description of the cluster sample included in the GCSL set of simulations.
Galaxy clusters $z_{\mathrm{cl}}$
Training + Validation Set ($4000+500$ images)
RX J2129++0005 (R2129) $0.234$
Abell 2744 (A2744) $0.308$
MACS J1931-2635 (M1931) $0.352$
MACS 1115++0129 (M1115) $0.352$
MACS J0416-2403 (M0416) $0.397$
MACS J1206-0847 (M1206) $0.439$
MACS J0329-0211 (M0329) $0.450$
RX J1347-1145 (R1347) $0.451$
MACS J2129-0741 (M2129) $0.587$
Test Set ($500$ images)
Abell S1063 (A1063) $0.348$
2.1 Euclid
The Euclid mission is a space-based survey by the European Space Agency that uses a telescope equipped with a $1.2\,\mathrm{m}$ mirror (Euclid Collaboration: Mellier et al. 2025). The Euclid Wide Survey (EWS; Euclid Collaboration: Scaramella et al. 2022) will map approximately $14\,000\,\deg^{2}$ of extragalactic sky with low zodiacal background and low Galactic extinction. The Euclid Deep Survey, instead, will cover about $50\,\deg^{2}$, achieving depths two magnitudes deeper than the wide survey. The first Euclid Quick Data Release (Q1; Euclid Quick Release Q1 2025) comprises $63.1\,\deg^{2}$ of the Euclid Deep Fields to nominal wide-survey depth, and contains about 30 million objects (Euclid Collaboration: Aussel et al. 2025).
[…]
2.2 HST2EUCLID
HST2EUCLID is a code developed by Euclid Collaboration: Bergamini et al. (2025b) to create simulated Euclid images in the $I_{\scriptscriptstyle\rm E}$, $Y_{\scriptscriptstyle\rm E}$, $J_{\scriptscriptstyle\rm E}$, and $H_{\scriptscriptstyle\rm E}$, bands starting from real HST observations. Although originally designed to produce mock Euclid observations of galaxy clusters, this tool can also be applied to any sufficiently deep HST image to create the corresponding Euclid-like observations, provided that the relevant HST filters overlap with the Euclid photometric bands that we want to simulate. Specifically, the HST/F606W and HST/F814W filters are combined to simulate the Euclid $I_{\scriptscriptstyle\rm E}$ band, while from the HST/F105W, HST/F125W, and HST/F160W filters we simulate the $Y_{\scriptscriptstyle\rm E}$, $J_{\scriptscriptstyle\rm E}$, and $H_{\scriptscriptstyle\rm E}$ bands. For further details on the methodology and implementation, we refer the reader to the original publication.
The HST2EUCLID simulation pipeline was validated through a series of three tests on the resulting Euclidised images. First, a check that the simulated images reach the expected depth of the EWS observations. Then, the distribution of galaxy sizes were measured in blank fields. Finally, the galaxy number counts detected in the Euclidised images were estimated to ensure consistency with expectations.
In our work, we exploited the Euclidised images in the four Euclid photometric bands of $10$ massive galaxy clusters (see Table 1), obtained with HST2EUCLID. We then simulated the GCSL events, and co-added them to the generated Euclidised image.
2.3 Lensing simulations
For the GCSL event simulations, we utilise deflection angle maps derived from high-precision SL models of $10$ galaxy clusters as presented by Caminha et al. (2019) and Bergamini et al. (2019, 2021, 2023a, 2023b, see Table 1) [^4] These lens models are publicly available at https://www.fe.infn.it/astro/lensing/. with the lenstool software (Kneib et al. 1996; Jullo et al. 2007; Jullo & Kneib 2009). These models utilise extensive spectroscopic data of multiple images to accurately represent both the large-scale mass distribution of the cluster and the sub-halo mass distribution (i.e. the cluster member galaxies). This detailed modelling captures the impact on the morphology, brightness, and occurrence of GCSL events.
Each cluster’s total mass distribution is represented by a parametric model of the overall lensing potential. This model incorporates the following components:
  • A cluster-scale component, consisting of DM halos and, when available, the smooth intra-cluster hot-gas mass from Chandra X-ray observations (Bonamigo et al. 2017, 2018).
  • A sub-halo mass component associated with cluster galaxies. The mass density profile of each sub-halo, including both dark and baryonic matter, is typically characterised by either a spherical or an elliptical singular dual-pseudo isothermal profile (Limousin et al. 2005; Elíasdóttir et al. 2007). This profile is further refined using measured stellar velocity dispersions from extensive samples of cluster member galaxies (Bergamini et al. 2021).
  • Any other massive objects in the galaxy cluster outskirts or along the line of sight.
[…]
lenstool (Kneib et al. 1996; Jullo et al. 2007; Jullo & Kneib 2009) reconstructs the cluster potential by minimising the offset between the observed positions of multiple images and those predicted by the model, given a specific set of model parameters. The reduced deflection angle, $\boldsymbol{\alpha}$, characterises the relation between the true position of the source $\boldsymbol{\beta}$ and its observed location $\boldsymbol{\theta}$ using the lens equation (Meneghetti 2021)
$\boldsymbol{\beta}=\boldsymbol{\theta}-\boldsymbol{\alpha}(\boldsymbol{\theta})\;.$ (1)
The simulation process is implemented using the Python library PyLensLib (Meneghetti 2021), and the key steps can be summarised as follows.
  1. 1.
    Given the mass distribution and redshift of the lens galaxy cluster and the redshift of the source, we numerically compute the deflection angle maps. Then, we derive the convergence and shear maps, these being the elements of the Jacobian matrix describing the image distortion, whose inverse matrix is known as the ‘magnification tensor’. Critical curves are then identified as the set of points where magnification diverges to infinity. An example of a tangential critical curve for the galaxy cluster Abell S1063 ($z_{\mathrm{cl}}=0.348$) and a source at redshift $z_{\mathrm{s}}=2.92$ is shown in red in Fig. 1 (left panel).
  2. 2.
    The primary critical line is mapped into the corresponding caustic line on the source plane (central panel in Fig. 1) using the lens equation (Eq. 1).
  3. 3.
    We simulate the lensing event by injecting a Sérsic surface brightness profile (Sérsic 1963, 1968), $I_{\mathrm{s}}(\boldsymbol{\beta})$, near the primary caustic line, within a buffer of width 0.5. We choose only the points within the buffer with a magnification factor $\mu$ greater than $50$. Since lens mapping conserves surface brightness, we can relate the observed and intrinsic surface brightnesses, $I(\boldsymbol{\theta})$ and $I_{\mathrm{s}}(\boldsymbol{\beta})$, using the lens equation $I(\boldsymbol{\theta})=I_{\mathrm{s}}(\boldsymbol{\beta})$; this allows us to directly reconstruct the source image in the lens plane through ray tracing (Meneghetti 2021).
  4. 4.
    Finally, the GCSL event is produced by convolving the simulated arcs with the Euclid PSF for each band, and then co-adding it to the HST2EUCLID base image in each filter (right panel in Fig. 1).
[…]
For the injected sources, we use a source spectral energy distribution (SED) based on a star-forming galaxy template from Kinney et al. (1996). Table 2 lists the Sérsic parameters along with their adopted value ranges. The Sérsic index, $n$, is sampled from a uniform distribution within $[1.0,2.0]$, which corresponds to typical starburst profiles of late-type galaxies. The axis ratio $q$ and position angle $\varphi$ are randomly drawn from uniform distributions in the ranges $[0.2,1.0]$ and $0,\pi]$, respectively. To closely reproduce the observed properties of the Sérsic sources, we follow the approach of Angora et al. (2023) and adopt a non-uniform sampling of the remaining parameters. For the source intrinsic magnitudes and redshifts we estimated the number counts in the $i$-band (that is, the number of galaxies per square degree per magnitude bin) from the COSMOS 2015 catalogue (Scoville et al. 2007; Laigle et al. 2016). Moreover, we complemented this with HST Deep Field North and South observations (Williams et al. 1996; Metcalfe et al. 2001) in F814W (Capak et al. 2007), which extends the galaxy counts to the faint end, down to $\mathrm{F814W}=29$. Then, we used the COSMOS photometric redshift catalogue to fit a redshift probability density function (PDF), i.e. $p(z\,|\Delta i)$, in six magnitude bins, using a function of the form $p(z\,|\Delta i)=Az^{2}\mathrm{e}^{-z/z_{0}}$ for $i\in[22,24)$ and $p(z\,|\Delta i)=Az^{2}\mathrm{e}^{-(z/z_{0})^{1/2}}$ for the other magnitude bins (Lombardi & Bertin 1999; Lombardi et al. 2005); see also Fig. 3 in Angora et al. 2023. We then use these PDFs to assign a source $I_{\scriptscriptstyle\rm E}$ magnitude to each background galaxy and, given this magnitude, a corresponding redshift.
For the injected sources, we retained only those with sampled magnitudes up to one magnitude fainter than the Euclid $I_{\scriptscriptstyle\rm E}$-band limiting magnitude, i.e. up to $I_{\scriptscriptstyle\rm E}=25.5$ (Euclid Collaboration: Scaramella et al. 2022). Following Meneghetti et al. (2022), we also enforce a lower limit for the source redshift of $z_{\mathrm{s}}=z_{\mathrm{cl}}+0.4$. Indeed, their study, which examined the lensing cross-section of the galaxy clusters used in this work, showed that the lensing cross-section increases substantially beyond zero for source redshifts roughly greater than $z_{\mathrm{cl}}+0.4$.
Finally, to assign an effective radius $r_{\mathrm{eff}}$ to the injected background galaxies, we use an empirical relation from Shibuya et al. (2015), describing how galaxy sizes evolve with redshift. This relation takes the form $r_{\mathrm{eff}}=B(1+z)^{\beta}$, based on galaxy size estimates in both optical and ultraviolet bands. However, since a comparison with the effective radii measured by Tortorelli et al. (2018) for low-redshift galaxies shows a significant overestimation, we apply this relation only for $z>1$, adopting a constant $r_{\mathrm{eff}}$ for $z\leq 1$ (see left panel of Fig. 4 in Angora et al. 2023).
Table 2: List of Sérsic parameters and their adopted value ranges for the injected sources.
Parameter Sampling strategy
Coordinate ($y_{\mathrm{s}}$, source plane) Sampled within a buffer
around the main caustic
Source magnitude ($i$) Sampled from PDF, $p(i)$
COSMOS + HST fields
Source redshift ($z_{\mathrm{s}}$) Sampled from PDF, $p(z|\Delta i)$
COSMOS
Effective radius ($r_{\mathrm{eff}}$) $r_{\mathrm{eff}}(z)=2.54$ kpc, $z\leq 1$
$r_{\mathrm{eff}}(z)=B(1+z)^{\beta}$, $z>1$
(Shibuya et al. 2015)
Sérsic index ($n$) Sampled from $\mathcal{U}[1.0,2.0]$
Axis ratio ($q$) Sampled from $\mathcal{U}[0.2,1.0]$
Position angle ($\varphi$) Sampled from $\mathcal{U}[0,\pi]$
2.4 Building the knowledge base
The described method enables the simulation of a large number of realistic GCSL events involving bright gravitational arcs. The cluster’s lens model is required to produce the knowledge base consisting of these lensing simulations. In addition, all galaxy clusters used in the training set already exhibit real cluster-scale bright gravitational arcs. The training set was generated as follows
  1. 1.
    Use the $10$ available highly-accurate lens models of the HST clusters, as done in Angora et al. (2023, see Table 1).
  2. 2.
    Generate Euclid-like images of these galaxy clusters using HST2EUCLID (Euclid Collaboration: Bergamini et al. 2025b).
  3. 3.
[…]
The combined training and validation set consists of $4500$ Euclid-like galaxy cluster images, each with dimensions of $120^{\prime\prime}\times 120^{\prime\prime}$. On these images, simulated gravitational lensed arcs (in addition to the existing ones) were added by randomly injecting Sérsic sources in the background, near the regions of higher magnification ($\mu\geq 50$). We kept the injected arcs only if their areas were larger than $400$ pixels; this area cut was implemented to focus the training specifically on the detection of giant and bright arcs, which are the primary targets of this work. This approach avoids the ambiguity inherent in smaller, fainter lensed arcs that could be confused by the NN with other objects, e.g. edge-on galaxies or stellar spikes, which could potentially degrade the network’s performance and increase the rate of false positives. This simulation process enables the direct extraction of bounding boxes and binary masks of the injected arcs. Conversely, the masks of pre-existing gravitational arcs were manually extracted for each cluster.
All the $500$ simulations corresponding to the last galaxy cluster in Table 1, namely A1063, were excluded from the training set. Instead, they were used exclusively as a test set for evaluating the performance of the NN.
3 Network and training
In this section, we describe our implementation of the Mask R-CNN for detecting bright gravitational lensing arcs in galaxy clusters. Our training procedure follows a supervised learning approach, utilising simulated SL events (and corresponding masks) on top of Euclidised images.
3.1 Implementation
Object detection and classification involve identifying objects within an image and assigning them to predefined classes. Semantic segmentation, on the other hand, labels each pixel of an image according to its class, without distinguishing between individual object instances. Instance segmentation combines these tasks by both detecting objects and generating a separate segmentation mask for each individual instance (Mueed Hafiz & Mohiuddin Bhat 2020). He et al. (2017) proposed the Mask R-CNN, an advanced framework for instance segmentation, extending the Faster R-CNN architecture (Girshick 2015; Ren et al. 2015). Mask R-CNN adds a branch for pixel-level segmentation and operates in two stages: first, generating object proposals and classifications, and then applying bounding boxes and segmentation masks to the image (see Fig. 2).
The Mask R-CNN architecture offers several key advantages over traditional CNNs when applied to lens detection tasks.
  • Instance segmentation. Unlike conventional CNNs that primarily perform image classification or basic object detection, Mask R-CNN generates detailed pixel-level segmentation masks for each detected object. It seamlessly integrates classification, bounding box regression, and segmentation tasks, whereas standard CNNs typically focus on a single task, such as whole-image classification.
  • Multi-object detection. Mask R-CNN is capable of detecting and segmenting multiple objects within a single image, even at different scales, while effectively distinguishing individual instances. In contrast, traditional CNNs often face challenges in identifying and separating overlapping objects.
  • Flexible input handling. Mask R-CNN can process images of varying sizes without the need to resize them to a fixed input dimension, unlike many standard CNN implementations. This flexibility helps preserve fine spatial details and prevents the loss of morphological information that can occur by interpolation during image resizing–an important advantage when working with high-resolution astrophysical data.
  • No need to produce the ‘negatives’. In the Mask R-CNN architecture every object that is not flagged as belonging to the ‘positive’ class (i.e. gravitational arcs) automatically pertains to the ‘negative’ one, removing in this way the need to create a fraction of the training set containing no arcs.
The code used in this work is based on the Python [^5] https://www.python.org/ implementation of the Mask R-CNN from torchvision (TorchVision 2016), built on the pytorch framework (Paszke et al. 2019; Ansel et al. 2024). The initial stage of the Mask R-CNN utilises a pre-trained CNN, known as the ‘backbone’, to extract feature maps from input images. The standard Mask R-CNN model (He et al. 2017, maskrcnn_resnet50_fpn [^6] https://pytorch.org/vision/stable/models/generated/torchvision.models.detection.maskrcnn_resnet50_fpn.html. in torchvision) utilises as backbone the ResNet-50 residual neural network (He et al. 2015). ResNets function as a feature extractor, with their initial layers capturing low-level features, such as edges and corners, and their deeper layers capturing high-level features via residual learning. ResNets are CNNs that employ ‘skip connections’, which allow the creation of deep architectures with many layers while avoiding the accuracy degradation issues commonly affecting deep NNs (He et al. 2015). This accuracy degradation happens because, as network depth increases, accuracy initially saturates and then swiftly declines (Bengio et al. 1994; He et al. 2017). The root cause of this problem lies in the backpropagation process, where continual multiplications by small weights diminish gradient sizes to the point of ineffectiveness; this issue is commonly known as the ‘vanishing gradient problem’ (Pascanu et al. 2012).
To handle objects of varying scales, features are extracted at different backbone stages and combined using a feature pyramid network (FPN; Lin et al. 2016), which integrates multiscale features by sharing information across hierarchical layers (Lakshmanan et al. 2021). The extracted feature maps are processed by the region proposal network (RPN; Ren et al. 2015), which generates object proposals using a sliding window approach that defines anchor boxes of varying sizes and aspect ratios at each spatial location on the feature map (Ren et al. 2015; Elgendy 2020). The RPN is trained to classify anchors as containing an object or not, while refining their bounding boxes, using a dedicated loss function to address both classification and regression tasks. In our implementation, we employ a set of anchors of sizes $\{32,64,128,256,512\}$ pixels, with aspect ratios $\{0.5,1,2\}$. The subsampling factor of ResNet-50, i.e. the ratio by which the spatial dimensions of an image are reduced as it passes through the NN, is $32$.
Non-maximum suppression (NMS; Girshick 2015) is subsequently applied to the proposed anchor boxes. This procedure begins by sorting all generated boxes according to their objectness scores, i.e. the confidence estimated by the network that a given region contains an object of interest. The algorithm then iteratively selects the box with the highest score and compares it against all remaining boxes using the intersection over union (IoU) metric,
$\mathrm{IoU}=\frac{|A\cap B|}{|A\cup B|}\;,$ (2)
where $A$ and $B$ denote two bounding boxes, the numerator represents the area of their intersection, while the denominator corresponds to the area of their union. To filter the RPN proposals, we adopt the default threshold of $\mathrm{IoU}_{\mathrm{thr}}^{\mathrm{RPN}}=0.7$ (He et al. 2017). Boxes that exhibit high IoU overlap with the selected box are suppressed, effectively removing redundant proposals. This process is repeated with the next highest-scoring unsuppressed box until all boxes have been processed or a predefined maximum number of proposals is reached. NMS thus helps to reduce computational load and enhances detection accuracy by retaining only the most promising, non-overlapping candidate regions.
The highest-ranked boxes, determined by their likelihood of containing an ‘object’, are selected as the region of interest (RoI) for the following stage of the pipeline. In our work, we categorise RoIs into two classes, ‘gravitational arcs’ and ‘background’; the latter, being trivial, will not be considered.
In the final stage of the Mask R-CNN, three tasks are performed simultaneously: a softmax classifier assigns probabilities for $K+1$ categories ($K=1$ in our case since we have only the ‘gravitational arcs’ class, while the +1+1 is for the background), a regression module refines bounding box coordinates, and a fully convolutional network generates pixel-wise binary masks within the RoI for semantic segmentation. The overall loss function of the Mask R-CNN is computed as the sum of these three contributions.
Figure 2: Mask R-CNN architecture. Adapted from Jung et al. (2019).
Transfer learning is a technique that enables NNs to leverage knowledge acquired from one task to enhance performance on a related but distinct task (refer to Tan et al. 2018 for an overview). This methodology capitalises on pre-existing weight values derived from one data set, employing them as the foundation for network initialisation for training the network on another data set. The advantages of this technique are twofold: it accelerates the training process, and it mitigates network overfitting. In our work, we utilise as a starting point the Mask R-CNN weights pre-trained on the Microsoft Common Objects in Context (MS COCO; Lin et al. 2014) data set. The MS COCO data set comprises approximately 328 000328\,000 images, categorised into $91$ classes of everyday objects. After initialising the pre-trained NN parameters, we made trainable all its $\approx$ 44 million parameters, including also the backbone ones.
3.2 Data preparation
Since the pytorch implementation of Mask R-CNN is designed for standard $3$-channel RGB images, an additional pre-processing step was required to retain as much information as possible. The third channel was then created by averaging the $J_{\scriptscriptstyle\rm E}$ and $H_{\scriptscriptstyle\rm E}$ Euclid bands (henceforth called the $JH_{\scriptscriptstyle\rm E}$ band). The three RGB channels are thus $JH_{\scriptscriptstyle\rm E}$, $Y_{\scriptscriptstyle\rm E}$, and $I_{\scriptscriptstyle\rm E}$, respectively.
Before feeding each image to the NN, we standardised the intensity of the pixels of each image. First, we clipped each band of the images independently at the $98{\mathrm{th}}$-percentile. Then, for each band of each image in the training set, we rescaled the pixel values using the $z$-score normalisation (Bishop & Bishop 2023), in order for the input image values to span a similar range of values, as
$R$ $=\left(JH_{\scriptscriptstyle\rm E}-\langle JH_{\scriptscriptstyle\rm E}\rangle\right)/\sigma_{JH_{\scriptscriptstyle\rm E}}\;,$ (3)
$G$ $=\left(Y_{\scriptscriptstyle\rm E}-\langle Y_{\scriptscriptstyle\rm E}\rangle\right)/\sigma_{Y_{\scriptscriptstyle\rm E}}\;,$ (4)
$B$ $=\left(I_{\scriptscriptstyle\rm E}-\langle I_{\scriptscriptstyle\rm E}\rangle\right)/\sigma_{I_{\scriptscriptstyle\rm E}}\;,$ (5)
where $\langle I_{\scriptscriptstyle\rm E}\rangle$ is the average value of the mean of the images in the $I_{\scriptscriptstyle\rm E}$-band, while $\sigma_{I_{\scriptscriptstyle\rm E}}$ is the average standard deviation, both computed over the whole training set (and similarly for the $Y_{\scriptscriptstyle\rm E}$- and $JH_{\scriptscriptstyle\rm E}$-bands). This allows a faster convergence of the NN, and prevents vanishing or exploding gradients (LeCun et al. 2012). This standardisation also ensures that the NN is not affected by the exposure time, detector gain, or the final normalisation of the pipeline-reduced images. The same data standardisation is then applied to real images during the inference phase.
We performed data augmentation on the training set, which makes the NN invariant with respect to the included transformations applied (Mikoajczyk & Grochowski 2018). In particular, we augmented the data by randomly applying for each image the following transformations, preserving classes, bounding boxes, and object masks:
  • horizontal flip, with $50\%$ probability;
  • vertical flip, with $50\%$ probability.
These basic augmentations replicate additional observational setups and conditions with minimal computational overhead, aiding the NN in better generalising its results.
3.3 Loss function
During training, each sampled RoI has a corresponding multitask loss, representing the three assignments of the Mask R-CNN (classification, location, and segmentation), expressed as $L_{\mathrm{RoI}}=L_{\mathrm{cls}}+L_{\mathrm{box}}+L_{\mathrm{mask}}$ (He et al. 2017), where:
  • the classification loss is $L_{\mathrm{cls}}=-\ln p_{u}$, where $p_{u}$ is the probability estimate from the NN for the true class $u$ and a discrete probability distribution $p=(p_{0},\dots,p_{K})$ is computed by a softmax on the output of a fully connected layer over $K+1$ classes for each RoI (Girshick 2015);
  • the bounding box loss $L_{\mathrm{box}}$ again follows the definition in Girshick (2015), using an $L1$ smooth loss of the offsets between predicted and real bounding boxes for each of the $K$ object classes;
  • for the mask branch, which contains $K$ binary masks of resolution $m\times m$ for each class ($m$ being the size of the RoI), a per-pixel sigmoid function is applied, leading to $L_{\mathrm{mask}}$ being the average binary cross-entropy loss (refer to He et al. 2017 for the details).
The final training loss $L$ is the combination of the classification, box, and mask losses from the final branches of the Mask R-CNN, along with the classification and box losses from the RPN (Lakshmanan et al. 2021).
3.4 Training
To train the NN, we used an improved version of the stochastic gradient descent algorithm (SGD; Goodfellow et al. 2016). SGD updates the model weights $\boldsymbol{\Theta}$ by minimising the loss function $L(\boldsymbol{\Theta})$, as described by
$\boldsymbol{\Theta}_{j+1}=\boldsymbol{\Theta}_{j}-\eta\frac{\partial}{\partial\boldsymbol{\Theta}_{j}}L(\boldsymbol{\Theta}_{j})\;,$ (6)
where $\eta$ represents the learning rate, a hyperparameter that is fine-tuned to avoid local minima and ensure convergence.
In all epochs, we train all the weights in all layers, using a learning rate scheduler, starting with a value of $\eta=10^{-3}$ and gradually decreasing it by a factor of $0.5$ every $20$ epochs. This approach facilitates deeper learning and more precise tuning of the weights, while at the same time reducing the risk of overfitting (Goodfellow et al. 2016). The SGD optimiser was initialised with a momentum value of $0.9$, and a weight decay coefficient of $5\times 10^{-4}$. Furthermore, to prevent overfitting we included the early stopping regularisation (Prechelt 1997; Raskutti et al. 2011).
We trained the NN for $100$ epochs in total, processing 4 0004\,000 training and $500$ validation $3\times 1199\times 1199\,\mathrm{pixel}$ images per epoch, in batches of $16$. We used a single NVIDIA Quadro RTX 6000 GPU, with $3840$ cores and $24\,\mathrm{GB}$ GDDR5 memory, to train on $4000$ simulated $3$-band FITS images. The training took approximately $10$ hours to complete (wall time). After this initial cost to train the NN, detection and inference on images of the same size can be performed in a fraction of a second.
4 Results
In this section, we present the results of the training, the performance of the NN, and the results of the application to real Euclid imaging data. We begin by testing the model’s performance using traditional ML metrics for classification tasks. Following this, we examine the outcomes of applying the trained Mask R-CNN to real Euclid Q1 observations, in order to assess its performance in discovering bright gravitational arcs in real galaxy cluster images.
4.1 Performance metrics
Two classes are present in our data set, the gravitational arcs (‘positive’ class) and everything else, i.e. the background and all other astronomical objects (‘negative’ class). We use the MS COCO Evaluation Metrics (Lin et al. 2014) to quantitatively evaluate the performance of the NN. In particular, to measure the classification performance on the test data set, we compute precision (or purity, $P$), recall (or completeness, $R$), and F1 score (their harmonic mean). Precision tells us how many of the positive predictions of the model are actually correct. Recall, on the other hand, is the ratio of correctly identified positive samples to the total number of actual positive samples in the data set. These metrics are defined as
$P$ $=\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FP}}\;,$ (7)
$R$ $=\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FN}}\;,$ (8)
$\mathrm{F1}$ $=2\frac{P\,R}{P+R}\;,$ (9)
where TP stands for true positive, the number of objects correctly classified as the positive class, TN for true negative, the number of objects correctly classified as the negative class, FP for false positive, the number of objects incorrectly labelled as belonging to the positive class, and FN for false negative, the number of objects incorrectly labelled as belonging to the negative class.
A detection is considered ‘positive’ if its classification confidence score exceeds a specified minimum threshold $p_{\mathrm{thr}}$, and such that its predicted bounding box overlaps with the ground truth one, giving an IoU larger than a threshold ($\mathrm{IoU}_{\mathrm{thr}}$). Therefore, TPs are identified when a detection has a confidence score above $p_{\mathrm{thr}}$ and can be matched to a ground truth object with $\mathrm{IoU}>\mathrm{IoU}_{\mathrm{thr}}$, FNs refer to ground truth objects that lack a corresponding detection, and FPs are detections with a high confidence score that do not correspond to any ground truth object.
Precision and recall are not particularly informative when considered separately; an object detector is considered effective only if its precision remains high as recall increases (Ivezić et al. 2020). Therefore, to measure the performance of the NN, we use the average precision (AP) score (Everingham et al. 2010), a widely adopted metric in the DL and computer vision communities, which has superseded the area under the receiver operating characteristic curve. The AP (for a specific class) is calculated as the area under the precision-recall (PPRR) curve, representing the precision averaged across uniformly distributed recall values from 0 to $1$. Generally, a higher precision at a given recall value indicates superior model detection performance. This is evidenced by a larger area under the PPRR curve; hence, higher AP values indicate better detection capabilities of the model. In this work, AP is computed for each image by averaging precision values at $101$ equally spaced recall levels, $R\in[0,0.01,\dots,1.0]$, namely as (Everingham et al. 2012)
$\mathrm{AP}=\frac{1}{101}\sum_{R\in[0,0.01,\dots,1.0]}P(R)\;,$ (10)
where $P(R)$ is the maximum precision within the interval $\Delta R$. We then calculate the average of the PPRR curves and mean AP scores across all test set images.
In the MS COCO evaluation metrics, the networks ability to distinguish and classify objects is assessed by computing AP in different ways. We use the standard AP score variants commonly used in the computer vision field. First, we repeat the computation of AP varying $\mathrm{IoU}_{\mathrm{thr}}$ from $0.50$ to $0.95$ in steps of $0.05$, resulting in $10$ sets of PPRR curves. The $10$ AP values derived from these $10$ sets are then averaged and denoted as $\mathrm{AP}_{@50:5:95}$, the primary COCO challenge metric. Then, $\mathrm{AP}_{50}$ (the Pascal VOC metric; Everingham et al. 2010) and $\mathrm{AP}_{75}$ are the AP metrics computed with $\mathrm{IoU}_{\mathrm{thr}}$ fixed at $0.5$ and $0.75$, respectively. Additionally, the AP for objects of different pixel areas is computed to evaluate the networks detection performance across various object sizes. For gravitational arcs detection, objects with an area smaller than $16^{2}$ pixels are denoted as $\mathrm{AP}_{\mathrm{S}}$, those with an area between $16^{2}$ and $32^{2}$ pixels are denoted as $\mathrm{AP}_{\mathrm{M}}$, and those larger than $32^{2}$ pixels are denoted as $\mathrm{AP}_{\mathrm{L}}$, calculated using the same IoU threshold range as that of $\mathrm{AP}_{@50:5:95}$ [^7] Notice that the values of $\mathrm{AP}_{\mathrm{S}}$, $\mathrm{AP}_{\mathrm{M}}$, and $\mathrm{AP}_{\mathrm{L}}$ used in our work are smaller with respect to those in the COCO Evaluation Metrics, since the pixel areas of gravitational arcs are generally smaller than typical COCO images.. The average recall (AR) instead is the maximum recall given a fixed number of detections per image ($100$ in our case), averaged over all the categories and IoU threshold values ranging from $0.50$ to $0.95$ in steps of $0.05$.
4.2 Performance on the test set
In Fig. 3 we show the training history of the total loss and its individual components as a function of the training epoch for both training and validation sets. The validation loss value stabilised after approximately $20$ iterations.
Figure 3: Training history of the total loss and all its individual components as a function of the training epoch, for the training (solid lines) and validation (dashed lines) data sets. The total loss (black lines) is the sum of the classification, box, and mask losses from the Mask R-CNN, along with the classification and box losses from the RPN.
As an example of the inference, we show in Fig. 4 the result of the application of the trained Mask R-CNN on a galaxy cluster belonging to the test set, and thus, not seen during the training phase. The green bounding boxes correspond to the ground truth, i.e. they enclose the gravitational arcs present in the galaxy cluster (both the real and simulated ones), while the red dashed boxes are the final proposed ‘positives’ by the NN. We can see from this example that the NN correctly recovers all the largest gravitational arcs, missing only the smaller ones and producing a few false positives.
Figure 4: Left: Euclidised $2^{\prime}\times 2^{\prime}$ RGB image ($R=JH_{\scriptscriptstyle\rm E}$, $G=Y_{\scriptscriptstyle\rm E}$, $B=I_{\scriptscriptstyle\rm E}$) of the galaxy cluster Abell 1063, belonging to the test set; the right-most arc is the one injected via SL simulation, while the others are real. Right: single channel $2^{\prime}\times 2^{\prime}$ Euclidised $I_{\scriptscriptstyle\rm E}$ image of the same cluster. The green boxes enclose the gravitational arcs (both real and simulated) present in the field, i.e. the ground truth, while the red dashed boxes are the ‘gravitational arcs’ found by the NN, having an object confidence score greater than $0.996$.
In Table 3, we present the COCO AP and AR metrics used to assess the performance of the NN on the test set, showing the models accuracy across various IoU thresholds and object sizes. These metrics offer insight into the strengths and limitations of the network. We observe that both AP and AR increase with arc size, with the best performance achieved in the large object category. This is expected, since larger arcs are more easily distinguishable from background structures and provide more pixels for the network to identify and segment. In contrast, smaller arcs are more affected by noise, and are often harder to resolve, given the resolution and depth of the images. This trend is also a direct consequence of the fact that our training set included only simulated arcs with an area larger than $400$ pixels, biasing the network toward high completeness and purity in the detection of large arcs. Furthermore, the comparison between $\mathrm{AP}_{50}$ ($54.3\%$) and $\mathrm{AP}_{75}$ ($20.1\%$) highlights that while the NN is relatively good at localising arcs with moderate overlap, it struggles to precisely match their shape and boundaries – a known challenge in segmenting faint, irregular sources.
[…]
$\mathrm{AP}_{@50:5:95}$ $\mathrm{AP}_{50}$ $\mathrm{AP}_{75}$ $\mathrm{AP}_{\mathrm{S}}$ $\mathrm{AP}_{\mathrm{M}}$ $\mathrm{AP}_{\mathrm{L}}$ $\mathrm{AR}_{\mathrm{S}}$ $\mathrm{AR}_{\mathrm{M}}$ $\mathrm{AR}_{\mathrm{L}}$
$21.8\%$ $54.3\%$ $20.1\%$ 1-1 $15.8\%$ $41.9\%$ 1-1 $18.1\%$ $57.9\%$
In Fig. 5 (a) we present the three PPRR curves for $\mathrm{IoU}_{\mathrm{thr}}=0.5,0.75$, and @5050:55:9595999@5050:55:$95$ means the value of the reference metric obtained by averaging it over IoU thresholds of $\mathrm{IoU}_{\mathrm{thr}}\in\{0.50,0.55,\dots,0.95\}$., each obtained by varying the object score classification threshold $p_{\mathrm{thr}}$. In Fig. 5 (b) we show the PPRR curve for $\mathrm{IoU}_{\mathrm{thr}}=0.5$, where the dots are colour coded by the score threshold $p_{\mathrm{thr}}$ used to evaluate them. Finally, in Fig. 5 (c) we show the precision and the recall as a function of the score threshold for $\mathrm{IoU}_{\mathrm{thr}}=0.5$. By choosing a fixed value of the object probability threshold, $p_{\mathrm{thr}}=0.996$, which can be tuned depending if we want to maximise precision over completeness, or vice versa, we can then compute the precision, recall, and F1 score of the NN on the test set, obtaining $P=75.9\%$, $R=58.0\%$, and $\mathrm{F1}=65.8\%$.
Figure 5: In panel (a) we show two sets of three PPRR curves for $\mathrm{IoU}_{\mathrm{thr}}=0.5,0.75$, @5050:55:$95$, each obtained by varying the object score threshold $p_{\mathrm{thr}}$. In panel (b) we show the PPRR curve for $\mathrm{IoU}_{\mathrm{thr}}=0.5$, where the dots are colour-coded by the score threshold $p_{\mathrm{thr}}$ used to evaluate them. In particular, we emphasise the PPRR pair associated to our choice of score threshold $p_{\mathrm{thr}}=0.996$. Note that the recall remains relatively high even at moderate precision, reflecting the network’s ability to recover most bright arcs while keeping the number of false positives manageable. Increasing the threshold shifts the operating point toward higher precision at the expense of recall. In panel (c) we show two precision and the recall curves as a function of the score threshold for $\mathrm{IoU}_{\mathrm{thr}}=0.5$.
We note that the test set is composed of images containing several positive examples by construction. In a realistic scenario, where most clusters do not host giant arcs, the effective precision would decrease due to the lower prevalence of true positives relative to false positives.
4.3 Inference on Euclid Q1 lens clusters
As a further test, we wanted to assess the performance of the NN on the $20$ galaxy clusters with lens probability $\mathcal{P}_{\mathrm{lens}}>0.90$ obtained from the Euclid Collaboration: Bergamini et al. (2025a) visual inspection. Euclid Collaboration: Bergamini et al. (2025a) presented the first catalogue of SL galaxy clusters identified in the Euclid Q1 observations, based on a systematic visual inspection of $1260$ richness-selected galaxy clusters from Wen & Han (2024), over an effective area of $4.4\,\deg^{2}$. They identified $83$ gravitational lenses with $\mathcal{P}_{\mathrm{lens}}>0.5$, including $14$ systems with $\mathcal{P}_{\mathrm{lens}}=1$ that exhibit secure SL features such as giant tangential and radial arcs, as well as multiple images.
In Fig. 6, we show in blue the distribution of the areas of the visually selected gravitational arcs [^10] These binary masks were manually drawn based on the bounding boxes surrounding the lensing events provided by the expert astronomers in Euclid Collaboration: Bergamini et al. (2025a).; in particular, we highlight the fact that the majority of them ($\approx 65\%$) have an area which is smaller than the minimum area used in the training set for the simulated arcs (i.e. $400$ pixels, indicated in the plot by the black dashed line). The red line instead shows the area distribution of the gravitational arcs correctly recovered by the NN. As expected, the majority of arcs with area larger than the training threshold are found ($12/18\approx 66\,\%$). Overall, by considering all the arcs found during the visual inspection in the 2020 $\mathcal{P}_{\mathrm{lens}}>0.90$ galaxy clusters as the ground truth for a further test set, the NN achieves a precision value of $P=19\%$, and recall of $R=42\%$. In Appendix A we show the predictions of the NN as bounding boxes over-plotted on the images of these 20 Euclid Q1 clusters, alongside the real arcs found during the visual inspection.
Figure 6: Distribution of area in pixels of the gravitational arcs found by expert astronomers in the 20 Euclid Q1 clusters with $\mathcal{P}_{\mathrm{lens}}>90\%$ (Euclid Collaboration: Bergamini et al. 2025a). The red dashed line shows the minimum value of the area of the arcs used as the ground truth for the training (400 pixels).
5 Discussion
In this paper, we have developed a DL methodology to autonomously identify cluster-scale SL gravitational arcs within Euclid multi-band imaging of galaxy clusters. The challenges of gravitational arc detection in wide-field surveys are well-suited for the Mask R-CNN architecture, an advanced computer vision framework offering several important advantages over traditional CNNs. First, it performs instance segmentation, generating detailed pixel-level masks for each detected object by integrating classification, bounding box regression, and segmentation into a unified framework. Standard CNNs, instead, are typically limited to single-task operations such as whole-image classification. In addition, Mask R-CNN can detect and segment multiple objects within a single image, even when they appear at different scales or overlap, effectively distinguishing individual instances, something conventional CNNs often struggle with. Another important advantage is interpretability: the pixel-level segmentation masks produced by the Mask R-CNN framework make it possible to visualise exactly which image regions the network has identified as arcs. This helps in understanding the origin of false positives, a task that is far more challenging with conventional CNNs that provide only image-level classifications. Finally, it can handle input images of varying sizes without requiring resizing to a fixed dimension, preserving structural details. In this context, our work represents one of the first applications of instance segmentation networks to the automated detection of cluster-scale gravitational arcs in real survey data, paving the way for future large-scale lensing searches in Euclid and other wide-field surveys.
The effectiveness of the network in detecting and classifying GCSLs within Euclid images is evaluated through several standard DL performance metrics. We also assess the capability of the NN to recover bright gravitational arcs in real Euclid Q1 observations, comparing the results of the inference with those obtained by human experts in a visual inspection (Euclid Collaboration: Bergamini et al. 2025a). Our analysis on the test set revealed that the Mask R-CNN framework performs well in arc detection, achieving a good level of precision-recall trade-off. Notably, the architectures ability to simultaneously localise, classify, and segment arcs streamlined the detection process and facilitated accurate identification in crowded fields such as those of galaxy clusters. The successful application of this framework to real Euclid Q1 images has validated its practical utility, with the majority of large arcs identified aligning with expert classifications.
Despite the promising results, certain limitations remain. Bright stars and elongated shapes occasionally mimicking arcs, produce false positives, highlighting the need for improved pre-processing of the images and more diversity in the training set. Increasing the prevalence of such challenging objects in the training set could help the network better distinguish true arcs from lens-like false positives. While the simulated data set effectively mimics real observations, the reliance on simulations may limit generalisability. As Euclid acquires more observational data, the performance of the Mask R-CNN is expected to improve for several reasons. With more real Euclid observations available, we can generate larger and more representative training data sets. This will help mitigate domain shift issues between simulated and real data, improving the generalization of the network. Moreover, new mock data sets can be generated with more realistic arc properties, taking into account improved lens models, observational systematics, and the full range of image conditions and artefacts in Euclid data. For instance, using real Euclid cluster images as the background onto which mock lensed sources are injected ensures that the simulations accurately reflect the properties of the actual observations, including realistic noise, PSF, photometric depth, and the presence of instrumental features such as diffraction spikes or saturated stars. Once a sufficiently large real data set is available, fine-tuning the existing Mask R-CNN model with actual Euclid images, rather than only relying on simulated arcs, will lead to better performance.
While the average precision and average recall test set results for large arcs ($\mathrm{AP}_{\mathrm{L}}$ and $\mathrm{AR}_{\mathrm{L}}$) are promising, the situation for medium- and small-sized ones reflected the inherent challenges in detecting these structures, which are often faint, blended with cluster galaxies, and affected by noise. In our work, we aimed particularly at detecting large bright gravitational arcs; therefore, the arcs injected in the training set reflected this goal. With continued improvements in training data sets, preprocessing, and model architecture, the detection of medium-sized arcs is expected to improve. However, small arcs will remain particularly challenging to detect due to their intrinsically low signal-to-noise ratios, and the higher abundance of small objects in the field that can mimic arc-like morphologies, compared to the relatively rarer large arc-like features that are not genuine arcs.
The Mask R-CNN framework was selected for this study due to its strong performance in instance segmentation and its ability to simultaneously classify, localise, and segment gravitational arcs. It is not necessarily the definitive or optimal solution, and there are several alternative architectures that could be explored in future work to further improve arc detection and segmentation, like U-Nets (Ronneberger et al. 2015) and YOLOs (Redmon et al. 2016); such networks, however, generally perform poorer on typical computer vision benchmarks (He et al. 2017).
Deploying the NN blindly across all Euclid tiles would inevitably lead to a significant number of false positives, which, while expected, would require extensive post-processing. The best strategy would be then to apply the network primarily to a selected subset of the richest galaxy clusters rather than across the entire survey tiles. This approach aims to balance minimising false positives with ensuring that genuine SL events are not missed. Moreover, while the DL models provide a fairly high level of accuracy in arc detection, a human validation step remains crucial to minimise false positives and ensure high purity in the final candidate sample. This step can be performed by expert astronomers visually inspecting the top-ranked candidates. A key advantage of our approach lies in its ability to significantly reduce the visual inspection effort required in large-scale surveys. In the Euclid Collaboration: Bergamini et al. (2025a) analysis of Euclid Q1 data, around $40$ experts visually examined approximately $1300$ candidate clusters, selected with a cut on richness from the Wen & Han (2024) catalogue. This process, while effective, is not scalable to future releases; indeed, Euclid Collaboration: Bergamini et al. (2025a) estimate that, with the same number of human experts, it would take more than $15$ years to visually inspect the entire EWS area.
This study further illustrates the potential of DL techniques in astrophysics, enabling the efficient and automated analysis of gravitational lensing phenomena in large-scale surveys. As current and next-generation surveys will produce unprecedentedly large data sets, advanced computer vision methodologies like this will be crucial in unlocking the full potential of these observations.
Acknowledgements.
We acknowledge financial support through grant 2020SKSTHZ. LB is indebted to the communities behind the multiple free, libre, and open-source software packages on which we all depend. MM was supported by INAF Grants The Big-Data era of cluster lensing” and ”Probing Dark Matter and Galaxy Formation in Galaxy Clusters through Strong Gravitational Lensing”, and ASI Grant n. 2024-10-HH.0 ”Attivit scientifiche per la missione Euclid fase E”. The research activities described in this paper have been co-funded by the European Union NextGeneration EU within PRIN 2022 project no. 20229YBSAN Globular clusters in cosmological simulations and in lensed fields: from their birth to the present epoch. The Euclid Consortium acknowledges the European Space Agency and a number of agencies and institutes that have supported the development of Euclid, in particular the Agenzia Spaziale Italiana, the Austrian Forschungsförderungsgesellschaft funded through BMIMI, the Belgian Science Policy, the Canadian Euclid Consortium, the Deutsches Zentrum für Luft- und Raumfahrt, the DTU Space and the Niels Bohr Institute in Denmark, the French Centre National d’Etudes Spatiales, the Fundação para a Ciência e a Tecnologia, the Hungarian Academy of Sciences, the Ministerio de Ciencia, Innovación y Universidades, the National Aeronautics and Space Administration, the National Astronomical Observatory of Japan, the Netherlandse Onderzoekschool Voor Astronomie, the Norwegian Space Agency, the Research Council of Finland, the Romanian Space Agency, the State Secretariat for Education, Research, and Innovation (SERI) at the Swiss Space Office (SSO), and the United Kingdom Space Agency. A complete and detailed list is available on the Euclid web site (www.euclid-ec.org). This work has made use of the Euclid Quick Release Q1 data from the Euclid mission of the European Space Agency (ESA), 2025, https://doi.org/10.57780/esa-2853f3b. This work used the following software packages: Python (Van Rossum & Drake 2009), PyTorch (Paszke et al. 2019; Ansel et al. 2024), torchvision (TorchVision 2016), NumPy (van der Walt et al. 2011; Harris et al. 2020), SciPy (Virtanen et al. 2020) scikit-learn (Pedregosa et al. 2011) Astropy (Astropy Collaboration et al. 2013; Price-Whelan et al. 2018), matplotlib (Hunter 2007), lenstool (Kneib et al. 1996; Jullo et al. 2007; Jullo & Kneib 2009), ds9 (Joye & Mandel 2003), Aladin (Bonnarel et al. 2000), topcat (Taylor 2005), git, bash (GNU 2007).
Appendix A Predictions on Euclid Q1 clusters
In this Appendix, we show the inference results of the trained Mask R-CNN on the 2020 $\mathcal{P}_{\mathrm{lens}}>0.90$ galaxy clusters from Euclid Collaboration: Bergamini et al. (2025a). The NN has been applied on a $2^{\prime}\times 2^{\prime}$ MER cutout on the nominal centre of the cluster, according to the catalogue. The red masks enclose the real arcs identified by expert astronomers; the dashed rectangles instead show the predicted ‘gravitational arcs’ by the NN, using the object score threshold $p_{\mathrm{thr}}=0.996$ employed for the test set (cf. Fig. 4). We note the presence of false positives, often associated with bright elongated galaxies, edge-on discs, or image artefacts that mimic arc-like morphologies, highlighting the need for further refinement of the training data or post-processing steps.
Figure 7: $2^{\prime}\times 2^{\prime}$ VIS cutouts of the 20 $\mathcal{P}_{\mathrm{lens}}>0.90$ galaxy clusters from Euclid Collaboration: Bergamini et al. (2025a). The red masks enclose the real arcs identified by expert astronomers; the dashed rectangles show the predictions of the NN.